The two acid ionization constants for oxalic acid, HOOCCOOH, are Ka1=5.6x10^-2 and Ka2=1.5x10^-4. What is the base dissociation constant, Kb, for OOCCOO2-(aq)?
1
Expert's answer
2020-06-11T10:30:56-0400

Kb2=Kw/Ka1=1.0 x 10^-14 / 5.6 x 10^-2=1.78e-13

Kb1=Kw/Ka2= 1.0 x 10^-14 / 1.5 x 10^-4=6.67e-11